What is the value of C
30 2?
a. 59
b. 60
c. 435
d. 870

Found 2 solutions by richard1234, jim_thompson5910:
Answer by richard1234(7193)

If you are referring to C(30,2), it is equal to 30*29/2!, or 15*29, or 435.

Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(30,2) = (30!)/(2!*(30-2)!)

C(30,2) = (30!)/(2!*28!)

C(30,2) = (30*29*28!)/(2!*28!)

C(30,2) = (30*29)/(2!)

C(30,2) = (30*29)/(2*1)

C(30,2) = (870)/(2)

C(30,2) = 435

So the answer is choice C